OpenStudy (anonymous):

is 2(rad2) the same thing as 2^(3/2)?

OpenStudy (anonymous):

\[2\sqrt{2} \ne 2^{\frac{3}{2}}\]

OpenStudy (anonymous):

actually it is true. think about it like this: \[\sqrt{2}\sqrt{2}\sqrt{2}\]

OpenStudy (anonymous):

then that would be \[\sqrt{2}^{3}\] which is equal to \[2^{3\over2}\]

OpenStudy (anonymous):

oooh, alright. thanks! that makes this problem way simpler!

OpenStudy (anonymous):

no problem! glad i could help! :)